OpenStudy (goformit100):

A water tank has steps inside it. A monkey is sitting on the topmost step (i.e., the first step). The water level is at the ninth step. (i) He jumps 3 steps down and then jumps back 2 steps up. In how many jumps will he reach the water level? (ii) After drinking water, he wants to go back. For this, he jumps 4 steps up and then jumps back 2 steps down in every move. In how many jumps will he reach back the top step?

OpenStudy (goformit100):

@Peter14

OpenStudy (anonymous):

a

OpenStudy (anonymous):

1st jump 1 + 3 = 4 2nd jump 4 + (−2) = 2 3rd jump 2 + 3 = 5 4th jump 5 + (−2) = 3 5th jump 3 + 3 = 6 6th jump 6 + (−2) = 4 7th jump 4 + 3 = 7 8th jump 7 + (−2) = 5 9th jump 5 + 3 = 8 10th jump 8 + (−2) = 6 11th jump, 6 + 3 = 9 Clearly, the monkey will be at water level i.e., 9th step after 11 jumps.